### v1

# A new bump in the night: Evidence of a new feature in the binary black hole mass distribution at $70~{}M_{\odot}$ from gravitational-wave observations

Abstract
We analyze the confident binary black hole (BBH) detections from the third Gravitational-Wave Transient Catalog (GWTC-3) with an alternative mass population model in order to capture features in the mass distribution beyond the Powerlaw + Peak model. We find that the peak of a second power law characterizes the $\sim 30-35~{}M_{\odot}$ bump, such that the data marginally prefers a mixture of two power laws for the mass distribution of binary components over a Powerlaw + Peak model with a Bayes Factor $\log_{10}\mathcal{B}$ of 0.1. This result may imply that the $\sim 30-35~{}M_{\odot}$ feature represents the onset of a second population of BBH mergers (e.g. from a dynamical formation channel) rather than a specific mass feature over a broader distribution. When an additional Gaussian bump is allowed within our power law mixture model, we find a new feature in the BH mass spectrum at $\sim 65-70~{}M_{\odot}$. This new feature may be consistent with hierarchical mergers, and constitute $\sim 2\%$ of the BBH population. This model also recovers a maximum mass of $58^{+30}_{-14}~{}M_{\odot}$ for the second power law, consistent with the onset of a pair-instability supernova mass gap.
gravitational waves, populations
1 Introduction
Since the first detection of a binary black hole (BBH) merger in gravitational waves (GW; Abbott et al. 2016a), the field of GW astronomy has entered the era of population studies. During the course of the first three observing runs conducted by LIGO and Virgo, 69 high significance BBH mergers have been detected (Abbott et al., 2023a), and hundreds are expected to be identified throughout the ongoing fourth LIGO-Virgo-KAGRA (LVK) observing run.
Multiple formation channels have been proposed to explain the observed population of compact objects beyond the standard isolated binaries scenario, including formation pathways requiring dynamical interactions in dense star clusters (e.g. Rodriguez et al. 2015, 2019), in the disks of Active Galactic Nuclei (AGN; e.g. McKernan et al. 2012; Antonini & Rasio 2016; McKernan et al. 2018) or in the center of dwarf galaxies (Conselice et al., 2020; Palmese & Conselice, 2021), black holes from metal-free Population III stars (Pop III; e.g. Kinugawa et al. 2014; Hartwig et al. 2016), and primordial black holes (e.g. Carr & Hawking 1974; Clesse & García-Bellido 2017; Tsai et al. 2021); see Mapelli (2021); Mandel & Farmer (2022) for reviews.
It is becoming increasingly clear that there is no single dominating channel responsible for the entire population, but rather a combination of formation mechanisms which can explain the various features and correlations observed in the data (Zevin et al., 2021; Cheng et al., 2023; Ray et al., 2024).
Multiple options for fitting the BBH mass distribution (and more broadly other population properties) have been proposed. The second and third GWTC fiducial analyses include a broken power law or a power law plus peak (Abbott et al., 2021a, b). Recent studies have also modeled the distribution with physics-motivated models (Baxter et al., 2021) or non-parametric approaches (Mandel et al., 2016; Abbott et al., 2021b; Ray et al., 2024; Magaña Hernandez & Ray, 2024; Farah et al., 2024a).
Motivated by the evidence for multiple BH populations, the presence of a peak at $\sim 30-35~{}M_{\odot}$ in the primary mass distribution, which is seemingly too narrow or at a location which is unrealistic to be explained by various formation channels (e.g. Golomb et al. 2023), the prediction for BBH from Pop III stars to peak around these values (Kinugawa et al., 2014), and the possibility of those being the seed of galaxies’ central black holes still residing in ultra-dwarf galaxies in the nearby Universe (Conselice et al., 2020; Palmese & Conselice, 2021), we explore whether the $\sim 30-35~{}M_{\odot}$ excess can be explained by a second power law rather than a peak over a broader power law, using the third GW Transient Catalog (GWTC-3; Abbott et al. 2023a).
Evidence for multiple BH populations includes the presence of a peak in the primary mass distribution at $\sim 30-35~{}M_{\odot}$, which is seemingly too narrow or too unrealistically located to to be explained by various formation channels Golomb et al. 2023). Furthermore, BBH from Pop III stars, which may be the seeds of of galaxies’ central black holes still residing in ultra-dwarf galaxies in the nearby Universe, are also predicted to peak around these values (Kinugawa et al., 2014; Conselice et al., 2020; Palmese & Conselice, 2021). Motivated by these theories, we explore whether the $\sim 30-35~{}M_{\odot}$ excess can be explained by a second power law rather than a peak over a broader power law by using the third GW Transient Catalog (GWTC-3) Abbott et al. 2023a).
Given that we are considering this possibility for a formation channel which is possibly dynamical, we also explore the presence of an additional high mass hierarchical population born from the peak of the power law.
We present our method in §2, show our results for GWTC-3 in §3, present validity tests using simulated GW events in §4, and discuss spin measurements in §5.
In §2 we present our method, in §3 we show our results on GWTC-3, in §4 we present validation tests using simulated GW events, and discuss spin measurements in §5. In §6 we discuss astrophysical implications of our findings, and in §8 we outline our summary and conclusions. Uncertainties throughout this work are at the 90% Credible Interval (CI).
2 Methods
We use hierarchical Bayesian population inference to simultaneously infer the parameters describing the mass and redshift distributions of BBH mergers with parametric models. We provide a brief summary below and refer the reader to Mandel et al. (2019); Thrane & Talbot (2019); Vitale et al. (2021) for an in depth description of the framework.
2.1 Hierarchical Bayesian Inference
The BBH observations provide us with an estimate of their primary mass $m_{1}$ and secondary mass $m_{2}$ components (in the source frame), as well as the measured redshift of the source $z$.
We can write the number density of BBH events as
$\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}=\frac{dV_{c}}{dz}\bigg{(}\frac{T_{\mathrm{obs}}}{1+z}\bigg{)}\mathcal{R}_{0}p(m_{1},m_{2},z|\Lambda)$ (1)
where $\Lambda$ denotes the hyperparameters describing the population distribution $p(m_{1},m_{2},z|\Lambda)$ of BBH masses and redshifts. Here, $dV_{c}/dz$ is the differential uniform-in-comoving volume element, $T_{\mathrm{obs}}$ is the total observation time, the factor of $1/(1+z)$ converts source-frame time to detector-frame time, and $\mathcal{R}_{0}$ is the local BBH merger rate at $z=0$. The BBH population is modeled through the normalized mass distribution $p(m_{1},m_{2},|\Lambda)$ as well as its redshift evolution $p(z|\Lambda)$, which we parameterize by a power law proportional to $(1+z)^{\kappa}$ as in previous studies (Fishbach et al., 2018; Abbott et al., 2023b).
Given a set of $N_{\mathrm{obs}}$ gravitational wave observations $\{d_{i}\}$, we can calculate the posterior on $\Lambda$ following Ref. Farr et al. (2019) and Mandel et al. (2019):
$\begin{split}p\left(\Lambda|\{d_{i}\}\right)&\propto p(\Lambda)e^{-\mathcal{R}_{0}\beta(\Lambda)}\prod_{i=1}^{N_{\mathrm{obs}}}\Bigg{[}\int\mathcal{L}\left(d_{i}|m_{1},m_{2},z\right)\\ &\times\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}dm_{1}dm_{2}dz\Bigg{]},\end{split}$ (2)
where $\mathcal{L}(d_{i}|m_{1},m_{2},z)$ is the single-event likelihood function for each event, and $\beta(\Lambda)$ is the detectable fraction of sources corresponding to a population determined by the population hyperparameters $\Lambda$.
Following Tiwari (2018) and Farr et al. (2019), we estimate $\beta(\Lambda)$ by assuming that it will follow a normal distribution (i.e. $\beta(\Lambda)\sim\mathcal{N}(\Lambda|\mu,\sigma)$), with $\mu$ as the importance sample estimate of $\beta$ and its associated uncertainty $\sigma^{2}=\mu^{2}/N_{\mathrm{eff}}$. Here $N_{\mathrm{eff}}$ is the number of effective samples which contribute to the estimate of $\beta$ given a population model described by $\Lambda$. As in Ref. Farr et al. (2019), we make sure that $N_{\mathrm{eff}}>4N_{\mathrm{obs}}$ to provide an unbiased estimate of $\beta$. Practically, we estimate the detectable fraction $\beta(\Lambda)$ by using the LVK’s injection campaign of BBH events simulated from a broad BBH population, injected into real detector data, and then searched for using the same analysis pipelines which found the events in GWTC-3 (Abbott et al., 2023b).
Finally, we marginalize over the local BBH merger rate $\mathcal{R}_{0}$ using a log-Uniform prior (Farr et al., 2019; Abbott et al., 2021a). We also approximate the integral over the individual event likelihoods in Eq. 2 with importance sampling over $N_{i}$ single-event posterior samples generated from single event inference analysis with the default event prior of $\pi(m_{1},m_{2},z,\chi_{\rm{eff}})\propto d_{L}^{2}(1+z)^{2}\frac{dd_{L}}{dz}$. For the $\chi_{\rm{eff}}$ inference in Sec. 5, we compute the default PE prior according to Callister (2021) using the gw-distributions library111publicly available at https://git.ligo.org/reed.essick/gw-distributions.
2.2 Model Comparison
Given the estimated hyper-parameter posterior distribution on $\Lambda$ for a given model $M$, we can compute the evidence for this model as:
[…]
If we have a competing model $N$, then the ratio of their evidences, the Bayes factor, provides us with a figure of merit to perform model comparison (Jeffreys, 1961). Here we use,
$\log_{10}\mathcal{B}^{M}_{N}=\log_{10}\left(\frac{Z_{M}}{Z_{N}}\right)$ (4)
where we have decided to work in log base 10 units for convenience. In general if $\log_{10}\mathcal{B}^{M}_{N}>0$, then model $M$ is favored over model $N$. Likewise, if $\log_{10}\mathcal{B}^{M}_{N}<0$, the opposite is true. If $\log_{10}\mathcal{B}^{M}_{N}\approx 0$, then it is inconclusive if one model is preferred over the other, and therefore both are equally good models at representing the underlying population given available data.
2.3 Population Models
Following Farah et al. (2024b), we model the population mass distribution as a symmetric distribution $p(m|\Lambda)$ on both the primary and secondary mass with a pairing function $f(m_{1},m_{2}|\Lambda)$. Here $\Lambda$ are the hyperparameters describing the mass models. That is, we consider the following family of models,
$p(m_{1},m_{2}|\Lambda)=p(m_{1}|\Lambda)p(m_{2}|\Lambda)f(m_{1},m_{2}|\Lambda).$ (5)
More specifically, we consider a pairing function which depends on the mass ratio $q=m_{2}/m_{1}$, or
$f(m_{1},m_{2}|\beta)=\left(\frac{m_{2}}{m_{1}}\right)^{\beta}.$ (6)
In this work, we consider the following family of models for $p(m|\Lambda)$ as a mixture of power laws $p_{{\rm{pl}}}$ and Gaussian $p_{{\rm{g}}}$ component distributions,
$p(m|\Lambda)=\sum_{i}f_{{\rm{pl}},i}p_{{\rm{pl}},i}(m|\Lambda)+\sum_{j}f_{{\rm{g}},j}p_{{\rm{g}},j}(m|\Lambda)$ (7)
where the mixture weights must add up to one or, $\sum_{i}f_{{\rm{pl}},i}+\sum_{j}f_{{\rm{g}},j}=1$.
More explicitly, we consider a mixture of two power laws and a single Gaussian component, and therefore we write, {widetext}
$p(m|\Lambda)=\sum_{i=1}^{2}f_{{\rm{pl}},i}S(m|m_{{\rm{min}},i},\delta m_{{\rm{min}},i})S(m|m_{{\rm{max}},i},\delta m_{{\rm{max}},i})p_{{\rm{pl}},i}(m|m_{{\rm{min}},i},m_{{\rm{max}},i},\alpha_{i})+f_{{\rm{g}}}p_{{\rm{g}}}(m|\mu_{m},\sigma_{m})$ (8)
where $\left(m_{{\rm{min}},i},m_{{\rm{max}},i},\alpha_{i}\right)$ correspond to the minimum mass, maximum mass, and power law slope for the $i$th power law component, respectively. $\mu_{m}$ and $\sigma_{m}$ represent the mean and standard deviation for the Gaussian component. The $S$ functions smooth each of the power law distribution components at both their lower and higher ends with smoothing scales $\delta m_{{\rm{X}},i}$ for $X=\{\rm{min},\rm{max}\}$. As before, the mixture weights add up to unity, i.e. $f_{{\rm{pl}},1}+f_{{\rm{pl}},2}+f_{{\rm{g}}}=1$.
With this in mind, we consider the following population models for the mass distribution of BBH mergers. First, if we set $f_{{\rm{pl}},2}=0$ we obtain the model with a single power law and a single Gaussian component, denoted by Powerlaw+Bump (PLB). This corresponds to the model described in detail in Farah et al. (2024b), however we do not refer to it as Powerlaw+Peak to avoid confusion with the model in Abbott et al. (2023a). Second, we set the Gaussian component mixture weight to zero, i.e. $f_{{\rm{g}}}=0$, and consider a model described by a mixture of two power laws, dubbed here as Powerlaw+Powerlaw (PLPL). Finally, we also consider the three-distribution mixture model described in Eq. 8 and denote it by Powerlaw+Powerlaw+Bump (PLPLB).
3 Results
In this section, we show the inferred population distributions for the masses of BBH mergers using the models we described in Section 2.3. Following the same GWTC-3 (Abbott et al., 2023c) population analysis choices (Abbott et al., 2023b), we only use the BBH events which pass a 1 per year False Alarm Rate (FAR) threshold. Therefore, we analyze a total of 69 confident BBH detections and exclude population outliers such as GW190814 from our analysis. For each event used in our analysis, we make use of the parameter estimation (PE) samples released by LVK (Abbott et al., 2023a). We also make use of the publicly available LVK GWTC-3 sensitivity injection campaign (Abbott et al., 2021b) to estimate the detection fraction $\beta(\Lambda)$ due to selection effects.
We show the main results of this paper in Fig. 1. We plot the marginal posterior distributions for both the primary and secondary masses for all models considered in this work. We also show the LVK GWTC-3 Powerlaw+Peak result for reference. We note that this distribution differs from PLB because it parameterizes $(m_{1},q)$ rather than $(m_{1},m_{2})$, and consequently sets different distribution shapes on both $m_{1}$ and $m_{2}$, with the peak present only in the primary mass distribution (Farah et al., 2024b).
Figure 1: Marginalized primary (top panel) and secondary (lower panel) mass distributions inferred with the PLB (pink), PLPL (blue) and PLB (green) models. For reference, we also show the GWTC-3 Powerlaw+Peak constraints in gray from Abbott et al. (2023b).
First, we note that the PLPL model from a qualitative perspective appears to be a good fit to the GWTC-3 data since the 90% confidence interval on the marginal mass distributions are consistent with Powerlaw+Peak. More quantitatively, we can compute the Bayes factor comparing the PLPL model to the PLB model to obtain an estimate of how well the two models represent the GWTC-3 population with respect to each other. We obtain $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}=+0.1$, and determine that the PLPL model describes the GWTC-3 observations similarly well to the PLB model, if not marginally better. In other words, the Gaussian bump at around $30-35~{}M_{\odot}$ could be described by a power law which is steeper than that recovered with the PLB model, turning on at a minimum mass of $m_{{\rm{min}},2}\sim 18~{}M_{\odot}$ and maximum mass $m_{{\rm{max}},2}\sim 45-100~{}M_{\odot}$ (at a different location depending on which model is chosen between PLPL and PLPLB), potentially changing the astrophysical interpretation of this population.
Second, we note that the more complex model, namely the PLPLB model, appears to find a feature at approximately $70~{}M_{\odot}$. In Fig. 2, we show the marginalized posterior distribution on $\{\mu_{m},\sigma_{m},f_{\rm{g}}\}$. Under the PLPLB model, we find consistency on the measured power law slope and minimum/maximum masses for the first power law compared to the PLPL model. As for the second power law, we find a consistent measurement for $m_{{\rm{min}},2}$ but find that its maximum mass shifts to a lower mass $m_{{\rm{max}},2}\sim~{}50M_{\odot}$ with a well defined mode compared to being preferentially larger than $\sim 70~{}M_{\odot}$. We find broad but consistent posteriors on $\alpha_{2}$ and the smoothing scales. The presence of the $\sim 70~{}M_{\odot}$ feature is robust with respect to the prior on the position $\mu_{m}$ of the Gaussian peak. For more details, we refer the reader to Appendix A for corner plots showing the posterior distributions of each model considered in this work.
[…]
Finally, if we extend the PLPLB model to have a second Gaussian component between $[20-50]~{}M_{\odot}$ (as in the PLB model), the feature at $\sim 30-35~{}M_{\odot}$ is still dominated by the second power law and the conclusions of this study are not changed.
Figure 2: Marginalized posterior distribution on the location $\mu_{m}$, width $\sigma_{m}$ and mixture weight $f_{\rm{g}}$ for the PLPLB model. We find the contribution to the overall population is 1-4%.
4 Validation
We validate our results using a large set of mock GW observations drawn from a known fiducial population model. Specifically, we want to test whether a PLPL or PLPLB model can spuriously fit the data given a BBH population which actually follows a PLB model. We draw simulated BBH events using the default LVK Powerlaw + Peak model (Talbot & Thrane, 2018) for the primary mass distribution. Mass ratios are drawn from a power law with slope $\beta$, and we allow the overall merger rate to evolve with redshift as a power law in $(1+z)^{\kappa}$. Our simulations are constructed to closely follow the inferred GWTC-3 population parameters from Abbott et al. (2023b), hence the fiducial choice for the hyper-parameter values for the simulated population described in Magaña Hernandez & Ray (2024). For reference, the marginalized component mass distributions for our simulated population are shown in Figure 3.
Subsequently, we assign observed values to detectable BBH systems, i.e. those which meet our selection criteria and calibrate our measurement uncertainty to what is expected for BBH mergers with the Advanced LIGO design sensitivity (Fishbach et al., 2018, 2020; Abbott et al., 2016b). We also simulate a corresponding set of detectable BBH systems from a broad fiducial population to take selection effects into account (Tiwari, 2018; Farr et al., 2019).
The goal of our validation study is to understand the distribution of $\log_{10}\mathcal{B}$ obtained on GWTC-3-sized catalogs. To do this, we use 69 mock BBH observations (drawn from our simulated catalog) and repeat this procedure so that we have 100 independent realizations in total. We then perform hierarchical population inference on each of our GWTC-3-sized catalogs with each of the population models considered in this work. We then compute the corresponding $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}$ and $\log_{10}\mathcal{B^{\textsc{PLPLB}}_{\textsc{PLB}}}$ factors for each catalog realization. In Fig. 4, we show the distribution of $\log_{10}\mathcal{B}$ for both the PLPL and PLPLB models (compared to PLB) and show the corresponding GWTC-3 results from Section 3 for reference.
Concerning the PLPL model, there are some simulated realizations which show a preference over the PLB model, i.e., $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}>0$ even when our simulated population was constructed with Powerlaw+Peak as the fiducial model. We find that $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}>0$ in 7% of our simulations. This shows that given the current size of the GWTC-3 catalog and its measurement uncertainty, we cannot confidently claim that PLPL is a preferred model over PLB. This is in agreement with the corresponding $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}$ for the GWTC-3 data, as can be seen from the overlap with the estimated distribution of Bayes factors.
More interestingly, we see that for the PLPLB model, given our assumptions on the simulation inputs, it is unlikely to reproduce a $\log_{10}\mathcal{B^{\textsc{PLPLB}}_{\textsc{PLB}}}$ factor consistent with the one measured with GWTC-3. We also note that for all the simulated realizations, we always find $\log_{10}\mathcal{B^{\textsc{PLPLB}}_{\textsc{PLB}}}<0$. This means that PLPLB is not a preferred representation of the underlying population if it is drawn from Powerlaw+Peak. Thus, our results and validation simulations potentially hint that a sub-population of BBH mergers at around $70~{}M_{\odot}$ is real, rather than a spurious feature due to the small number statistics and the specific representation we observed with GWTC-3.
As a secondary check, in Fig. 3, we plot the the 90% confidence interval constructed from the median component mass distribution reconstructions for each of the simulated GWTC-3-sized catalogs using the mass models considered in this work. We first note that PLB, PLPL, and PLPLB are all able to recover the underlying Powerlaw+Peak population.
For the PLPLB model, we are most interested in determining how often and with which amplitude the Gaussian component fits a non-existent feature in the range of $60-80~{}M_{\odot}$. This is done to understand how sensitive our model is to the potentially small number statistics for the GWTC-3 catalog realizations and how, for some realizations, we might get clusters of events in the $60-80~{}M_{\odot}$ mass range when no sub-population is present. We see that this can happen, however, conditional on our simulation assumptions we see that we cannot reproduce our PLPLB GWTC-3 results, e.g. the amplitude of the Gaussian feature that we measure is a lot stronger than what we see in our simulations.
To quantify the fraction of realizations which can reproduce our results due to clustering of events around $70~{}M_{\odot}$ when no such sub-population is present, in Fig. 5 we show the marginalized posterior distribution on $f_{\rm{g}}$ for both our simulations and GWTC-3. We find that for GWTC-3, $f_{\rm g}>0.01$ is at 95% confidence. For our simulations, only 3% of the realizations have a similar significance of $f_{\rm{g}}>0.01$ at 95% confidence.
Figure 3: Primary (top panel) and secondary (lower panel) component mass distributions (90% confidence interval) from the median mass distribution reconstructions for each of the 100 simulated GWTC-3-sized catalogs described in Section 4 for the PLB (pink), PLPL (blue) and PLB (green) models. For reference, we show the simulated GWTC-3 Powerlaw+Peak like population as the solid black line.
Figure 4: In pink (green) we show the distribution of $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}$ ($\log_{10}\mathcal{B^{\textsc{PLPLB}}_{\textsc{PLB}}}$) for the simulated GWTC-3-sized catalogs described in Sec. 4. The vertical lines show the corresponding GWTC-3 results.
Figure 5: Marginalized posterior distributions (gray) for each of the 100 realizations of our GWTC-3-like simulations on the mixture model weight $f_{\rm{g}}$ for the Gaussian component in the PLPLB model. In green, we show the measured posterior on $f_{\rm{g}}$ with GWTC-3 data.
5 Does spin tell us anything?
Apart from masses and redshifts, BBH observations provide us with information about each black hole spin orientation and magnitude through the effective inspiral spin parameter $\chi_{\rm{eff}}$ defined as,
$\chi_{\rm{eff}}=\frac{m_{1}\chi_{1,z}+m_{2}\chi_{2,z}}{m_{1}+m_{2}}$ (9)
where $\chi_{1,z}$ and $\chi_{2,z}$ are the dimensionless spin vector components (projected along the angular momentum axis of the binary), for the primary and secondary BH respectively. In general, $\chi_{\rm{eff}}$ is a measure for the alignment ($\chi_{\rm{eff}}>0$) or misalignment ($\chi_{\rm{eff}}<0$) of individual BH spins.
From a population point of view, the distribution of $\chi_{\rm{eff}}$ can provide hints as to the origin and formation channels of BBH mergers. A positively skewed $\chi_{\rm{eff}}$ distribution is consistent with the majority of mergers having positively aligned spins, hinting at formation in the field (Gerosa et al., 2018; Zaldarriaga et al., 2018; Belczynski et al., 2020), whereas a distribution which is symmetric about zero, hints at dynamically formed binaries due to the spin tilts being isotropically distributed at binary assembly (Rodriguez et al., 2015; Antonini & Rasio, 2016; Gerosa & Berti, 2017).
Thus to probe the origin of the sub-populations discussed in Sec. 3 for the PLPLB model, we consider extended models $p(m,\chi_{\rm{eff}}|\Lambda)$ which fit the population distribution of $\chi_{\rm{eff}}$ for each component in our mixture models. We model the $\chi_{\rm{eff}}$ distribution as a truncated Gaussian distribution in the range $\chi_{\rm{eff}}\in[-1,1]$ for each sub-population, that is, we have three distinct $\chi_{\rm{eff}}$ distributions with population parameters $\{\mu_{\chi,1},\sigma_{\chi,1},\mu_{\chi,2},\sigma_{\chi,2},\mu_{\chi,\rm{g}},\sigma_{\chi,\rm{g}}\}$ describing the location and width of the $\chi_{\rm{eff}}$ distributions for the first and second power law, as well as the Gaussian bump, respectively.
In Fig. 6, we show the posterior distribution on the $\chi_{\rm{eff}}$ population distribution parameters and in Fig. 7, we show the reconstructed primary mass and $\chi_{\rm{eff}}$ population distributions for each of the components for the PLPLB model. From our results, we find that $\mu_{\chi,1}>0$ at $>99.8$% confidence for the first power law (PL1) component, whereas for the second power law PL2 we find $\mu_{\chi,2}$ and $\mu_{\chi,\rm{g}}$ to be constrained to the within the ranges $[-0.06,0.26]$ and $[-0.92,0.41]$ respectively at 90% confidence. We find this population to be broad but skewed towards negative $\chi_{\rm{eff}}$. However, more data in this region of parameter space is needed to make definitive conclusions regarding the origin of BHs in this sub-population.
Consequently, when we look at the reconstructed component distributions in primary mass and $\chi_{\rm{eff}}$, we see that PL1 is slightly skewed towards positive $\chi_{\rm{eff}}$ and is well constrained within the range $[-0.2,0.2]$. As for PL2, we find that it is consistent with a distribution symmetric about zero but has some preference for positive $\chi_{\rm{eff}}$. This is not surprising, as the mass range for $m_{1}$ for PL2 is in the range $20-60~{}M_{\odot}$ and may include higher mass objects formed through dynamical interaction as well as some lower mass objects from field binaries. More interestingly, for the sub-population at $70~{}M_{\odot}$, we find a broad and symmetric distribution on $\chi_{\rm{eff}}$.
Figure 6: Marginalized posterior distribution on the locations and standard deviations for the Gaussian $\chi_{\rm{eff}}$ population distributions for each mixture component of the PLPLB model.
Figure 7: Reconstructed $m_{1}$ and $\chi_{\rm{eff}}$ population distributions for each of the components in the PLPLB model.
6 Astrophysical Implications
7 Low-mass $\sim 10-20~{}M_{\odot}$ population
We start by discussing the first power law, PL1, which appears to be consistent with the expectations for isolated binaries (e.g. van Son et al. 2022), and in agreement with findings from previous works searching for isolated binaries from this sub-population using GWTC-3 (Godfrey et al., 2024; Li et al., 2024; Ray et al., 2024). This population is likely to drop more steeply with mass than what observed in (Abbott et al., 2021b) for the broader power law, with a maximum mass buried under PL2. Its effective spin distribution is likely (at 90% CI) centered at values larger than zero, therefore supporting marginally aligned (with the binary angular momentum) spins and/or slowly spinning black holes. Aligned spins are typically predicted within the isolated binary formation channel which is in agreement with our observations. On this other hand, assuming this origin, the relatively low values may imply efficient angular momentum transport between the envelope and the core of the progenitor star leading to black holes that are slowly spinning (Bavera et al., 2020; Belczynski et al., 2020), and supernova kicks introducing misalignment between spins and binary angular momentum (Kalogera, 1996). Note that other formation channels are also expected to produce BBHs within this population, so we cannot assume that the entirety of this population (comprising $\sim 77\%$ of the BBH population according to the PLPLB model) originates in the field.
7.1 Intermediate-mass $\sim 20-40~{}M_{\odot}$ population
Concerning the PL2 population, we are interpreting the $\sim 30~{}M_{\odot}$ peak as the onset of a second BH population, rather than a special feature on its own. The PLPLB model favors a maximum mass at $\sim 45-50~{}M_{\odot}$, while the PLB reaches $\sim 100~{}M_{\odot}$. Of course in the latter case, the model needs to also fit for the high mass mergers with component masses at $\sim 60~{}M_{\odot}$, so the second PL needs a higher maximum mass. Interestingly, the maximum mass for the PLPLB is $m_{\rm{max},2}=58^{+30}_{-14}~{}M_{\odot}$, and is consistent with the expectation for the onset of the pair–instability supernova (PISN) mass gap (Farmer et al., 2019), in which case it is possible that the PL2 population also originates from isolated binaries, or is made of first generation dynamically-formed binaries.
For PL2 from PLPLB, we find a smoothing kernel of $\sim 20~{}M_{\odot}$, revealing a less narrow feature than in the PLB model (found at $32^{+1.9}_{-2.0}~{}M_{\odot}$ with width $3^{+4.0}_{-1.9}~{}M_{\odot}$ in the PLB model) which may be more easily explained by popular formation channels, which typically do not predict such a narrow feature at $\sim 30-35~{}M_{\odot}$.
Various formation channels may be able to reproduce a second $>20~{}M_{\odot}$ mass population in the mass distribution which is marginally better represented by a second power law rather than a Gaussian bump. Note that this is different from a broken power law as the amplitude of the second power law is free to vary, such that this second population does not necessarily need to be connected to the one peaking at $\sim 10~{}M_{\odot}$. Broekgaarden et al. (2022); van Son et al. (2022, 2023); Cheng et al. (2023) show that binary stellar population models can produce various mass distribution features, including a $\sim 30-35~{}M_{\odot}$ peak. For instance, depending on whether the binary underwent a common envelope phase (CE), one could distinguish between an isolated binary CE subchannel and a stable Roche-lobe overflow (RLOF) subchannel (van Son et al., 2022), which may explain the emergence of two power laws below and above $\sim 30~{}M_{\odot}$. Additionally, dense stellar clusters, including nuclear and globular clusters, can predict a sub-population peaking around this mass (Mapelli et al., 2022; Cheng et al., 2023).
The spin distribution is relatively inconclusive with respect to the origin of PL2. Although consistent with being centered at zero, it is still marginally skewed towards aligned spins - possibly because it originates from dynamical channels but is contaminated by objects which belong to the lower mass population, or because this sub-population also originates in the field. The larger width of the distribution compared to PL1 is compelling for a dynamical origin. Biscoveanu et al. (2022); Heinzel et al. (2024b) finds a correlation between the width of the effective spin distribution and redshift. We argue that this may be related to what we observe here. If we allow for the different sub-populations to vary with redshift independently of each other, we find that at higher redshift the lower mass population rate decreases (as also found by Rinaldi et al. 2024) while the PL2 rate increases. While such redshift evolution may not be physical due to the limited detector horizon especially at lower masses (Heinzel et al., 2024a), this finding may reconcile our and previous works’ observations that the second sub-population drives the spreading of the spin width at higher redshift. However, note that uncertainties on spins for higher mass systems are typically larger than those from PL1, so it is challenging to draw definitive conclusions. Perhaps the most compelling evidence that PL2 may be composed of isolated binaries or dynamically formed first generation binaries is the fact that the PLPLB model PL2 has a maximum mass consistent with PISN mass gap. An involve study exploring the redshift and $\chi_{\rm{eff}}$ dependence on the mass distribution is ongoing.
We note that Godfrey et al. (2024) and (Li et al., 2024) also find evidence for 2 distinct sub-populations. Specifically, Godfrey et al. (2024) finds that the higher mass population which we identify here with PL2 shares characteristics with the likely isolated binary population peaking at $\sim 10~{}M_{\odot}$. Given all of the arguments and previous works outlined, we argue that PL1 is composed of isolated binaries, while PL2 may mostly arise from 1g+1g binaries from dynamical formation.
7.2 High-mass $\sim 60-80~{}M_{\odot}$ population
Although it may be tempting to identify the feature at $\sim 70~{}M_{\odot}$ as the pulsational PISN (PPISN) build-up (Talbot & Thrane, 2018), the bump is located at a mass too high to be at the lower edge of the PISN mass gap, which is expected to be around $45~{}M_{\odot}$ and likely below $56~{}M_{\odot}$ considering reasonable variations in metallicity and uncertainties in the CO reaction rates (Farmer et al., 2019). For the high mass bump, we find a broad distribution on $\chi_{\rm{eff}}$. Although with large uncertainties given the current statistics, this population supports the presence of effective spins $<-0.3$, which are expected to exist for hierarchical mergers (Baibhav et al., 2020; Fishbach et al., 2022), but are hardly explained by isolated binary formation. Given these arguments, we argue that a plausible explanation for this feature is from hierarchical mergers of BHs from the $\sim 30-35~{}M_{\odot}$ bump. Because we find that the fraction of BBHs in the bump is 1-4% at 90% confidence, assuming these are from 2g BHs in dense star clusters, it is likely that these BHs are born spinning ($\chi_{\rm birth}>0$), but not enough to wash out this feature completely ($\chi_{\rm birth}<0.5$), according to the predictions of Rodriguez et al. 2019. A high mass distribution similar to what found here is also predicted by Kremer et al. (2020), with the bump above $\sim 60~{}M_{\odot}$ originating from 2g hierarchical mergers and BHs born from stellar collision.
8 Conclusions
In this paper we have explored alternative mass models for the BBH mass population. In particular, we construct a mixture model of power law and Gaussian distributions and fit the confident GWTC-3 BBH observations to date. We find that a mixture of two power law distributions is equally preferred by the GWTC-3 data compared to a power law plus Gaussian peak. The first power law is likely comprised by a large fraction of field binaries, which is consistent with previous findings, while we argue that the second power law includes a significant fraction of dynamically assembled first generation compact object binaries. However, our simulations show that a mixture of two power laws may model an underlying PLB population similarly well to what we find on the data for $\sim 7\%$ of our representations, and claim that more data is needed to robustly assess whether the PLPL model is indeed preferred over the PLPLB one.
We also find that adding a Gaussian component above $40~{}M_{\odot}$ robustly finds a sub-population located around $65-70~{}M_{\odot}$. It is possible that by allowing a second power law to exist with a steeper slope than the first PL, one enables the detection of a sub-population that was otherwise buried under a single, flatter power law. This sub-population may arise from second generation hierarchical mergers, forming from black holes in the secondary peak of the distribution at $\sim 35~{}M_{\odot}$. Our simulations show that a similar feature cannot be recovered by chance coincidence when the underlying population follows a PLB. However, more data is required to confidently claim that the PLPLB model is preferred over other models. We expect that the increasing number of BBH events detected through O4 and beyond will provide the observations required to determine whether the $70~{}M_{\odot}$ sub-population is actually present in the astrophysical BH mass distribution. If confirmed, such sub-population would a smoking gun about the dynamical origin of high-mass LVK BBHs, and may represent a new valuable feature to improve cosmological parameters constraints through spectral standard siren measurements (Farr et al., 2019; Ezquiaga & Holz, 2022).
9 Acknowledgements
The authors would like to thank Anarya Ray, Katelyn Breivik, Caitlin Rose and Yeajin Kim for useful comments. IMH is supported by a McWilliams postdoctoral fellowship at Carnegie Mellon University. This material is based upon work supported by the National Aeronautics and Space Administration under Grant No. 22-LPS22-0025. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
Appendix A Full Posterior distributions
Figure 8: Corner plot showing the full posterior distribution for the PLB model.
Figure 9: Corner plot showing the full posterior distribution for the PLPL model.
Figure 10: Corner plot showing the full posterior distribution for the PLPLB model.

### v2

@@
 We analyze the confident binary black hole (BBH) detections from the third Gravitational-Wave Transient Catalog (GWTC-3) with an alternative mass population model in order to capture features in the mass distribution beyond the Powerlaw + Peak model.
-We find that the peak of a second power law characterizes the $\sim 30-35~{}M_{\odot}$ bump, such that the data marginally prefers a mixture of two power laws for the mass distribution of binary components over a Powerlaw + Peak model with a Bayes Factor $\log_{10}\mathcal{B}$ of 0.1.
+We find that the peak of a second power law characterizes the $\sim 30-35~{}M_{\odot}$ bump, such that the data marginally prefers a mixture of two power laws for the mass distribution of binary components over a Powerlaw + Peak model with a Bayes Factor $\log_{10}\mathcal{B}$ of 0.24.
 This result may imply that the $\sim 30-35~{}M_{\odot}$ feature represents the onset of a second population of BBH mergers (e.g.
 from a dynamical formation channel) rather than a specific mass feature over a broader distribution.
-When an additional Gaussian bump is allowed within our power law mixture model, we find a new feature in the BH mass spectrum at $\sim 65-70~{}M_{\odot}$.
-This new feature may be consistent with hierarchical mergers, and constitute $\sim 2\%$ of the BBH population.
-This model also recovers a maximum mass of $58^{+30}_{-14}~{}M_{\odot}$ for the second power law, consistent with the onset of a pair-instability supernova mass gap.
-gravitational waves, populations
-1 Introduction
-Since the first detection of a binary black hole (BBH) merger in gravitational waves (GW; Abbott et al.
-2016a), the field of GW astronomy has entered the era of population studies.
+When an additional Gaussian bump is allowed within our power law mixture model, we find a new feature in the BH mass spectrum at $\sim 65-70~{}M_{\odot}$ ($\log_{10}\mathcal{B}$ = 0.29 compared to Powerlaw + Peak).
+This new feature may be consistent with hierarchical mergers, and constitute $\sim 3\%$ of the BBH population.
+This model also recovers a maximum mass of $58^{+32}_{-14}~{}M_{\odot}$ for the second power law, consistent with the onset of a pair-instability supernova mass gap.
+gravitational-waves, populations
+I Introduction
+Since the first detection of a binary black hole (BBH) merger in gravitational waves (GW) Abbott et al.
+(2016a), the field of GW astronomy has entered the era of population studies.
 During the course of the first three observing runs conducted by LIGO and Virgo, 69 high significance BBH mergers have been detected (Abbott et al., 2023a), and hundreds are expected to be identified throughout the ongoing fourth LIGO-Virgo-KAGRA (LVK) observing run.
-Multiple formation channels have been proposed to explain the observed population of compact objects beyond the standard isolated binaries scenario, including formation pathways requiring dynamical interactions in dense star clusters (e.g.
-Rodriguez et al.
-2015, 2019), in the disks of Active Galactic Nuclei (AGN; e.g.
-McKernan et al.
-2012; Antonini & Rasio 2016; McKernan et al.
-2018) or in the center of dwarf galaxies (Conselice et al., 2020; Palmese & Conselice, 2021), black holes from metal-free Population III stars (Pop III; e.g.
-Kinugawa et al.
-2014; Hartwig et al.
-2016), and primordial black holes (e.g.
-Carr & Hawking 1974; Clesse & García-Bellido 2017; Tsai et al.
-2021); see Mapelli (2021); Mandel & Farmer (2022) for reviews.
+Multiple formation channels have been proposed to explain the observed population of compact objects beyond the standard isolated binaries scenario, including formation pathways requiring dynamical interactions in dense star clusters (Rodriguez et al., 2015, 2019), in the disks of Active Galactic Nuclei (AGN) (McKernan et al., 2012; Antonini and Rasio, 2016; McKernan et al., 2018) or in the center of dwarf galaxies (Conselice et al., 2020; Palmese and Conselice, 2021), black holes from metal-free Population III stars (Pop III) (Kinugawa et al., 2014; Hartwig et al., 2016), and primordial black holes (Carr and Hawking, 1974; Clesse and García-Bellido, 2017; Tsai et al., 2021).
+For in depth reviews, see (Mapelli, 2021; Mandel and Farmer, 2022).
 It is becoming increasingly clear that there is no single dominating channel responsible for the entire population, but rather a combination of formation mechanisms which can explain the various features and correlations observed in the data (Zevin et al., 2021; Cheng et al., 2023; Ray et al., 2024).
 Multiple options for fitting the BBH mass distribution (and more broadly other population properties) have been proposed.
-The second and third GWTC fiducial analyses include a broken power law or a power law plus peak (Abbott et al., 2021a, b).
-Recent studies have also modeled the distribution with physics-motivated models (Baxter et al., 2021) or non-parametric approaches (Mandel et al., 2016; Abbott et al., 2021b; Ray et al., 2024; Magaña Hernandez & Ray, 2024; Farah et al., 2024a).
-Motivated by the evidence for multiple BH populations, the presence of a peak at $\sim 30-35~{}M_{\odot}$ in the primary mass distribution, which is seemingly too narrow or at a location which is unrealistic to be explained by various formation channels (e.g.
-Golomb et al.
-2023), the prediction for BBH from Pop III stars to peak around these values (Kinugawa et al., 2014), and the possibility of those being the seed of galaxies’ central black holes still residing in ultra-dwarf galaxies in the nearby Universe (Conselice et al., 2020; Palmese & Conselice, 2021), we explore whether the $\sim 30-35~{}M_{\odot}$ excess can be explained by a second power law rather than a peak over a broader power law, using the third GW Transient Catalog (GWTC-3; Abbott et al.
-2023a).
-Evidence for multiple BH populations includes the presence of a peak in the primary mass distribution at $\sim 30-35~{}M_{\odot}$, which is seemingly too narrow or too unrealistically located to to be explained by various formation channels Golomb et al.
-2023).
-Furthermore, BBH from Pop III stars, which may be the seeds of of galaxies’ central black holes still residing in ultra-dwarf galaxies in the nearby Universe, are also predicted to peak around these values (Kinugawa et al., 2014; Conselice et al., 2020; Palmese & Conselice, 2021).
-Motivated by these theories, we explore whether the $\sim 30-35~{}M_{\odot}$ excess can be explained by a second power law rather than a peak over a broader power law by using the third GW Transient Catalog (GWTC-3) Abbott et al.
-2023a).
-Given that we are considering this possibility for a formation channel which is possibly dynamical, we also explore the presence of an additional high mass hierarchical population born from the peak of the power law.
-We present our method in §2, show our results for GWTC-3 in §3, present validity tests using simulated GW events in §4, and discuss spin measurements in §5.
-In §2 we present our method, in §3 we show our results on GWTC-3, in §4 we present validation tests using simulated GW events, and discuss spin measurements in §5.
-In §6 we discuss astrophysical implications of our findings, and in §8 we outline our summary and conclusions.
+The second and third Gravitational-Wave Transient Catalog (GWTC) fiducial analyses include a broken power law or a power law plus peak (Abbott et al., 2021a, b).
+Recent studies have also modeled the distribution with physics-motivated models (Baxter et al., 2021) or non-parametric approaches (Mandel et al., 2016; Abbott et al., 2021b; Edelman et al., 2023; Ray et al., 2024; Magaña Hernandez and Ray, 2024; Farah et al., 2024a).
+Evidence for multiple BH populations includes the presence of a peak in the primary mass distribution at $\sim 30-35~{}M_{\odot}$, which is seemingly too narrow or too unrealistically located to to be explained by various formation channels (Briel et al., 2023; Golomb et al., 2024; Hendriks et al., 2023).
+Furthermore, BBH from Pop III stars, which may be the seeds of of galaxies’ central black holes still residing in ultra-dwarf galaxies in the nearby Universe, are also predicted to peak around these values (Kinugawa et al., 2014; Conselice et al., 2020; Palmese and Conselice, 2021).
+Motivated by these theories, we explore whether the $\sim 30-35~{}M_{\odot}$ excess can be explained by a second power law rather than a peak over a broader power law by using the confidently detected BBH observations from the GWTC-3 catalog  (Abbott et al., 2023a)).
+Given that we are considering a population model for a formation channel which is possibly of dynamical origin (see Gerosa and Fishbach (2021) for a review), we also explore the presence of an additional high mass hierarchical merger population (Abbott et al., 2020a; Kimball et al., 2020, 2021) that could originate from the merger of two black holes each formed from the merger of another two compact objects in the $30-35~{}M_{\odot}$ peak of the power law.
+In §II we present our method, in §III we show our results on GWTC-3, in §IV we present validation tests using simulated GW events, and discuss spin measurements in §V.
+In §VI we discuss astrophysical implications of our findings, and in §VII we outline our summary and conclusions.
 Uncertainties throughout this work are at the 90% Credible Interval (CI).
-2 Methods
+II Methods
 We use hierarchical Bayesian population inference to simultaneously infer the parameters describing the mass and redshift distributions of BBH mergers with parametric models.
-We provide a brief summary below and refer the reader to Mandel et al.
-(2019); Thrane & Talbot (2019); Vitale et al.
-(2021) for an in depth description of the framework.
-2.1 Hierarchical Bayesian Inference
+We provide a brief summary below and we refer the reader to review articles for an in depth description of the framework (Mandel et al., 2019; Thrane and Talbot, 2019; Vitale et al., 2021).
+II.1 Hierarchical Bayesian Inference
 The BBH observations provide us with an estimate of their primary mass $m_{1}$ and secondary mass $m_{2}$ components (in the source frame), as well as the measured redshift of the source $z$.
 We can write the number density of BBH events as
-$\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}=\frac{dV_{c}}{dz}\bigg{(}\frac{T_{\mathrm{obs}}}{1+z}\bigg{)}\mathcal{R}_{0}p(m_{1},m_{2},z|\Lambda)$ (1)
-where $\Lambda$ denotes the hyperparameters describing the population distribution $p(m_{1},m_{2},z|\Lambda)$ of BBH masses and redshifts.
+$\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}\propto\frac{dV_{c}}{dz}\bigg{(}\frac{T_{\mathrm{obs}}}{1+z}\bigg{)}\mathcal{R}_{0}(1+z)^{\kappa}p(m_{1},m_{2}|\Lambda)$ (1)
+where $\Lambda$ denotes the hyperparameters describing the population distribution $p(m_{1},m_{2}|\Lambda)$ of BBH masses.
 Here, $dV_{c}/dz$ is the differential uniform-in-comoving volume element, $T_{\mathrm{obs}}$ is the total observation time, the factor of $1/(1+z)$ converts source-frame time to detector-frame time, and $\mathcal{R}_{0}$ is the local BBH merger rate at $z=0$.
-The BBH population is modeled through the normalized mass distribution $p(m_{1},m_{2},|\Lambda)$ as well as its redshift evolution $p(z|\Lambda)$, which we parameterize by a power law proportional to $(1+z)^{\kappa}$ as in previous studies (Fishbach et al., 2018; Abbott et al., 2023b).
-Given a set of $N_{\mathrm{obs}}$ gravitational wave observations $\{d_{i}\}$, we can calculate the posterior on $\Lambda$ following Ref. Farr et al. (2019) and Mandel et al. (2019):
-$\begin{split}p\left(\Lambda|\{d_{i}\}\right)&\propto p(\Lambda)e^{-\mathcal{R}_{0}\beta(\Lambda)}\prod_{i=1}^{N_{\mathrm{obs}}}\Bigg{[}\int\mathcal{L}\left(d_{i}|m_{1},m_{2},z\right)\\ &\times\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}dm_{1}dm_{2}dz\Bigg{]},\end{split}$ (2)
+The BBH population is modeled through the normalized mass distribution $p(m_{1},m_{2},|\Lambda)$ as well as its redshift evolution, which we parameterize by a power law proportional to $(1+z)^{\kappa}$ as in previous studies (Fishbach et al., 2018; Abbott et al., 2023b).
+Given a set of $N_{\mathrm{obs}}$ gravitational wave observations $\{d_{i}\}$, we can calculate the posterior on $\Lambda$ following (Farr et al., 2019) and (Mandel et al., 2019):
+$\begin{split}p\left(\Lambda|\{d_{i}\}\right)&\propto\frac{p(\Lambda)}{\beta(\Lambda)^{N_{\mathrm{obs}}}}\prod_{i=1}^{N_{\mathrm{obs}}}\Bigg{[}\int\mathcal{L}\left(d_{i}|m_{1},m_{2},z\right)\\ &\times\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}dm_{1}dm_{2}dz\Bigg{]},\end{split}$ (2)
 where $\mathcal{L}(d_{i}|m_{1},m_{2},z)$ is the single-event likelihood function for each event, and $\beta(\Lambda)$ is the detectable fraction of sources corresponding to a population determined by the population hyperparameters $\Lambda$.
@@
 We also approximate the integral over the individual event likelihoods in Eq.
-2 with importance sampling over $N_{i}$ single-event posterior samples generated from single event inference analysis with the default event prior of $\pi(m_{1},m_{2},z,\chi_{\rm{eff}})\propto d_{L}^{2}(1+z)^{2}\frac{dd_{L}}{dz}$.
-For the $\chi_{\rm{eff}}$ inference in Sec.
-5, we compute the default PE prior according to Callister (2021) using the gw-distributions library111publicly available at https://git.ligo.org/reed.essick/gw-distributions.
-2.2 Model Comparison
+2 with importance sampling over $N_{i}$ single-event posterior samples generated by the LVK from single event parameter estimation (PE) analyses with prior choices corresponding to $\pi(m_{1},m_{2},z)\propto d_{L}^{2}(1+z)^{2}\frac{dd_{L}}{dz}$ Abbott et al.
+(2023b).
+In Sec.
+V, we analyze additional spin information for each merger encoded in their $\chi_{\rm{eff}}$ measurement (a mass-weighted, aligned, effective spin, see Sec.
+V for its definition).
+Therefore, we compute the default PE prior $\pi(m_{1},m_{2},z,\chi_{\rm{eff}})$ according to Callister (2021) using the gw-distributions library111publicly available at https://git.ligo.org/reed.essick/gw-distributions.
+II.2 Model Comparison
 Given the estimated hyper-parameter posterior distribution on $\Lambda$ for a given model $M$, we can compute the evidence for this model as:
@@
 If $\log_{10}\mathcal{B}^{M}_{N}\approx 0$, then it is inconclusive if one model is preferred over the other, and therefore both are equally good models at representing the underlying population given available data.
-2.3 Population Models
+II.3 Population Models
 Following Farah et al.
@@
 That is, we consider the following family of models,
-$p(m_{1},m_{2}|\Lambda)=p(m_{1}|\Lambda)p(m_{2}|\Lambda)f(m_{1},m_{2}|\Lambda).$ (5)
+$p(m_{1},m_{2}|\Lambda)\propto p(m_{1}|\Lambda)p(m_{2}|\Lambda)f(m_{1},m_{2}|\Lambda).$ (5)
 More specifically, we consider a pairing function which depends on the mass ratio $q=m_{2}/m_{1}$, or
 $f(m_{1},m_{2}|\beta)=\left(\frac{m_{2}}{m_{1}}\right)^{\beta}.$ (6)
-In this work, we consider the following family of models for $p(m|\Lambda)$ as a mixture of power laws $p_{{\rm{pl}}}$ and Gaussian $p_{{\rm{g}}}$ component distributions,
+In this work, we consider the following family of models for $p(m|\Lambda)$ as a mixture of power laws $p_{{\rm{pl}}}$ and Gaussian $p_{{\rm{g}}}$ component probability distributions,
 $p(m|\Lambda)=\sum_{i}f_{{\rm{pl}},i}p_{{\rm{pl}},i}(m|\Lambda)+\sum_{j}f_{{\rm{g}},j}p_{{\rm{g}},j}(m|\Lambda)$ (7)
 where the mixture weights must add up to one or, $\sum_{i}f_{{\rm{pl}},i}+\sum_{j}f_{{\rm{g}},j}=1$.
-More explicitly, we consider a mixture of two power laws and a single Gaussian component, and therefore we write, {widetext}
+More explicitly, we consider a mixture of two power laws and a single Gaussian component, and therefore we write,
 $p(m|\Lambda)=\sum_{i=1}^{2}f_{{\rm{pl}},i}S(m|m_{{\rm{min}},i},\delta m_{{\rm{min}},i})S(m|m_{{\rm{max}},i},\delta m_{{\rm{max}},i})p_{{\rm{pl}},i}(m|m_{{\rm{min}},i},m_{{\rm{max}},i},\alpha_{i})+f_{{\rm{g}}}p_{{\rm{g}}}(m|\mu_{m},\sigma_{m})$ (8)
@@
 8 and denote it by Powerlaw+Powerlaw+Bump (PLPLB).
-3 Results
-In this section, we show the inferred population distributions for the masses of BBH mergers using the models we described in Section 2.3.
+III Results
+In this section, we show the inferred population distributions for the masses of BBH mergers using the models we described in Section II.3.
 Following the same GWTC-3 (Abbott et al., 2023c) population analysis choices (Abbott et al., 2023b), we only use the BBH events which pass a 1 per year False Alarm Rate (FAR) threshold.
-Therefore, we analyze a total of 69 confident BBH detections and exclude population outliers such as GW190814 from our analysis.
+Therefore, we analyze a total of 69 confident BBH detections and exclude population outliers such as GW190814 (Abbott et al., 2020b) from our analysis.
 For each event used in our analysis, we make use of the parameter estimation (PE) samples released by LVK (Abbott et al., 2023a).
@@
 More quantitatively, we can compute the Bayes factor comparing the PLPL model to the PLB model to obtain an estimate of how well the two models represent the GWTC-3 population with respect to each other.
-We obtain $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}=+0.1$, and determine that the PLPL model describes the GWTC-3 observations similarly well to the PLB model, if not marginally better.
+We obtain $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}=+0.24$, and determine that the PLPL model describes the GWTC-3 observations similarly well to the PLB model, if not marginally better.
 In other words, the Gaussian bump at around $30-35~{}M_{\odot}$ could be described by a power law which is steeper than that recovered with the PLB model, turning on at a minimum mass of $m_{{\rm{min}},2}\sim 18~{}M_{\odot}$ and maximum mass $m_{{\rm{max}},2}\sim 45-100~{}M_{\odot}$ (at a different location depending on which model is chosen between PLPL and PLPLB), potentially changing the astrophysical interpretation of this population.
@@
 […]
+[…]
 Finally, if we extend the PLPLB model to have a second Gaussian component between $[20-50]~{}M_{\odot}$ (as in the PLB model), the feature at $\sim 30-35~{}M_{\odot}$ is still dominated by the second power law and the conclusions of this study are not changed.
@@
 We find the contribution to the overall population is 1-4%.
-4 Validation
+[…]
+IV Validation
 We validate our results using a large set of mock GW observations drawn from a known fiducial population model.
 Specifically, we want to test whether a PLPL or PLPLB model can spuriously fit the data given a BBH population which actually follows a PLB model.
-We draw simulated BBH events using the default LVK Powerlaw + Peak model (Talbot & Thrane, 2018) for the primary mass distribution.
+We draw simulated BBH events using the default LVK Powerlaw + Peak model (Talbot and Thrane, 2018) for the primary mass distribution.
 Mass ratios are drawn from a power law with slope $\beta$, and we allow the overall merger rate to evolve with redshift as a power law in $(1+z)^{\kappa}$.
 Our simulations are constructed to closely follow the inferred GWTC-3 population parameters from Abbott et al.
-(2023b), hence the fiducial choice for the hyper-parameter values for the simulated population described in Magaña Hernandez & Ray (2024).
+(2023b), hence the fiducial choice for the hyper-parameter values for the simulated population described in Magaña Hernandez and Ray (2024).
 For reference, the marginalized component mass distributions for our simulated population are shown in Figure 4.
@@
 In Fig.
-4, we show the distribution of $\log_{10}\mathcal{B}$ for both the PLPL and PLPLB models (compared to PLB) and show the corresponding GWTC-3 results from Section 3 for reference.
+5, we show the distribution of $\log_{10}\mathcal{B}$ for both the PLPL and PLPLB models (compared to PLB) and show the corresponding GWTC-3 results from Section III for reference.
 Concerning the PLPL model, there are some simulated realizations which show a preference over the PLB model, i.e., $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}>0$ even when our simulated population was constructed with Powerlaw+Peak as the fiducial model.
@@
 As a secondary check, in Fig.
-3, we plot the the 90% confidence interval constructed from the median component mass distribution reconstructions for each of the simulated GWTC-3-sized catalogs using the mass models considered in this work.
+4, we plot the the 90% confidence interval constructed from the median component mass distribution reconstructions for each of the simulated GWTC-3-sized catalogs using the mass models considered in this work.
 We first note that PLB, PLPL, and PLPLB are all able to recover the underlying Powerlaw+Peak population.
@@
 6 we show the marginalized posterior distribution on $f_{\rm{g}}$ for both our simulations and GWTC-3.
-We find that for GWTC-3, $f_{\rm g}>0.01$ is at 95% confidence.
-For our simulations, only 3% of the realizations have a similar significance of $f_{\rm{g}}>0.01$ at 95% confidence.
-Figure 3: Primary (top panel) and secondary (lower panel) component mass distributions (90% confidence interval) from the median mass distribution reconstructions for each of the 100 simulated GWTC-3-sized catalogs described in Section 4 for the PLB (pink), PLPL (blue) and PLB (green) models.
+We find that for GWTC-3, $f_{\rm g}>0.01$ is at 96% confidence.
+For our simulations, only 1 of the realizations have a similar significance of $f_{\rm{g}}>0.01$ at 96% confidence.
+Figure 4: Primary (top panel) and secondary (lower panel) component mass distributions (90% confidence interval) from the median mass distribution reconstructions for each of the 100 simulated GWTC-3-sized catalogs described in Section IV for the PLB (pink), PLPL (blue) and PLB (green) models.
 For reference, we show the simulated GWTC-3 Powerlaw+Peak like population as the solid black line.
 Figure 5: In pink (green) we show the distribution of $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}$ ($\log_{10}\mathcal{B^{\textsc{PLPLB}}_{\textsc{PLB}}}$) for the simulated GWTC-3-sized catalogs described in Sec.
-4.
+IV.
 The vertical lines show the corresponding GWTC-3 results.
@@
 In green, we show the measured posterior on $f_{\rm{g}}$ with GWTC-3 data.
-5 Does spin tell us anything?
+V Does spin tell us anything?
 Apart from masses and redshifts, BBH observations provide us with information about each black hole spin orientation and magnitude through the effective inspiral spin parameter $\chi_{\rm{eff}}$ defined as,
@@
 From a population point of view, the distribution of $\chi_{\rm{eff}}$ can provide hints as to the origin and formation channels of BBH mergers.
-A positively skewed $\chi_{\rm{eff}}$ distribution is consistent with the majority of mergers having positively aligned spins, hinting at formation in the field (Gerosa et al., 2018; Zaldarriaga et al., 2018; Belczynski et al., 2020), whereas a distribution which is symmetric about zero, hints at dynamically formed binaries due to the spin tilts being isotropically distributed at binary assembly (Rodriguez et al., 2015; Antonini & Rasio, 2016; Gerosa & Berti, 2017).
+A positively skewed $\chi_{\rm{eff}}$ distribution is consistent with the majority of mergers having positively aligned spins, hinting at formation in the field (Gerosa et al., 2018; Zaldarriaga et al., 2018; Belczynski et al., 2020), whereas a distribution which is symmetric about zero, hints at dynamically formed binaries due to the spin tilts being isotropically distributed at binary assembly (Rodriguez et al., 2015; Antonini and Rasio, 2016; Gerosa and Berti, 2017).
 Thus to probe the origin of the sub-populations discussed in Sec.
-3 for the PLPLB model, we consider extended models $p(m,\chi_{\rm{eff}}|\Lambda)$ which fit the population distribution of $\chi_{\rm{eff}}$ for each component in our mixture models.
+III for the PLPLB model, we consider extended models $p(m,\chi_{\rm{eff}}|\Lambda)$ which fit the population distribution of $\chi_{\rm{eff}}$ for each component in our mixture models.
 We model the $\chi_{\rm{eff}}$ distribution as a truncated Gaussian distribution in the range $\chi_{\rm{eff}}\in[-1,1]$ for each sub-population, that is, we have three distinct $\chi_{\rm{eff}}$ distributions with population parameters $\{\mu_{\chi,1},\sigma_{\chi,1},\mu_{\chi,2},\sigma_{\chi,2},\mu_{\chi,\rm{g}},\sigma_{\chi,\rm{g}}\}$ describing the location and width of the $\chi_{\rm{eff}}$ distributions for the first and second power law, as well as the Gaussian bump, respectively.
 In Fig.
-6, we show the posterior distribution on the $\chi_{\rm{eff}}$ population distribution parameters and in Fig.
-7, we show the reconstructed primary mass and $\chi_{\rm{eff}}$ population distributions for each of the components for the PLPLB model.
-From our results, we find that $\mu_{\chi,1}>0$ at $>99.8$% confidence for the first power law (PL1) component, whereas for the second power law PL2 we find $\mu_{\chi,2}$ and $\mu_{\chi,\rm{g}}$ to be constrained to the within the ranges $[-0.06,0.26]$ and $[-0.92,0.41]$ respectively at 90% confidence.
+7, we show the posterior distribution on the $\chi_{\rm{eff}}$ population distribution parameters.
+In Fig.
+8, we show the reconstructed primary and secondary mass population distributions for each of the components for the PLPLB model.
+Similarly, in Fig.
+9 we show the reconstructed $\chi_{\rm{eff}}$ population distribution for each of the components for the PLPLB model.
+From our results, we find that $\mu_{\chi,1}>0$ at $>99.8$% confidence for the first power law (PL1) component and to be constrained within the range $[0.02,0.08]$ at 90% confidence.
+Whereas for the second power law (PL2) and Gaussian Bump we find $\mu_{\chi,2}$ and $\mu_{\chi,\rm{g}}$ to be constrained to the within the ranges $[-0.08,0.15]$ and $[-0.87,0.52]$ respectively at 90% confidence.
 We find this population to be broad but skewed towards negative $\chi_{\rm{eff}}$.
@@
 Figure 7: Marginalized posterior distribution on the locations and standard deviations for the Gaussian $\chi_{\rm{eff}}$ population distributions for each mixture component of the PLPLB model.
-Figure 7: Reconstructed $m_{1}$ and $\chi_{\rm{eff}}$ population distributions for each of the components in the PLPLB model.
-6 Astrophysical Implications
-7 Low-mass $\sim 10-20~{}M_{\odot}$ population
-We start by discussing the first power law, PL1, which appears to be consistent with the expectations for isolated binaries (e.g.
-van Son et al.
-2022), and in agreement with findings from previous works searching for isolated binaries from this sub-population using GWTC-3 (Godfrey et al., 2024; Li et al., 2024; Ray et al., 2024).
+Figure 8: Reconstructed $m_{1}$ and $m_{2}$ population distributions for each of the components in the PLPLB model.
+Figure 9: Reconstructed $\chi_{\rm{eff}}$ population distributions for each of the components in the PLPLB model.
+We also show the $\text{CDF}(\chi_{\rm{eff}})$ to more easily demonstrate the symmetry (about $\chi_{\rm{eff}}=0$) for each sub-population.
+VI Astrophysical Implications
+VI.1 Low-mass $\sim 10-20~{}M_{\odot}$ population
+We start by discussing the first power law, PL1, which appears to be consistent with the expectations for isolated binaries (van Son et al., 2022), and in agreement with findings from previous works searching for isolated binaries from this sub-population using GWTC-3 (Godfrey et al., 2024; Li et al., 2024; Ray et al., 2024).
 This population is likely to drop more steeply with mass than what observed in (Abbott et al., 2021b) for the broader power law, with a maximum mass buried under PL2.
@@
 On this other hand, assuming this origin, the relatively low values may imply efficient angular momentum transport between the envelope and the core of the progenitor star leading to black holes that are slowly spinning (Bavera et al., 2020; Belczynski et al., 2020), and supernova kicks introducing misalignment between spins and binary angular momentum (Kalogera, 1996).
-Note that other formation channels are also expected to produce BBHs within this population, so we cannot assume that the entirety of this population (comprising $\sim 77\%$ of the BBH population according to the PLPLB model) originates in the field.
-7.1 Intermediate-mass $\sim 20-40~{}M_{\odot}$ population
+Note that other formation channels are also expected to produce BBHs within this population, so we cannot assume that the entirety of this population (comprising $\sim 76\%$ of the BBH population according to the PLPLB model) originates in the field.
+VI.2 Intermediate-mass $\sim 20-40~{}M_{\odot}$ population
 Concerning the PL2 population, we are interpreting the $\sim 30~{}M_{\odot}$ peak as the onset of a second BH population, rather than a special feature on its own.
@@
 Of course in the latter case, the model needs to also fit for the high mass mergers with component masses at $\sim 60~{}M_{\odot}$, so the second PL needs a higher maximum mass.
-Interestingly, the maximum mass for the PLPLB is $m_{\rm{max},2}=58^{+30}_{-14}~{}M_{\odot}$, and is consistent with the expectation for the onset of the pair–instability supernova (PISN) mass gap (Farmer et al., 2019), in which case it is possible that the PL2 population also originates from isolated binaries, or is made of first generation dynamically-formed binaries.
-For PL2 from PLPLB, we find a smoothing kernel of $\sim 20~{}M_{\odot}$, revealing a less narrow feature than in the PLB model (found at $32^{+1.9}_{-2.0}~{}M_{\odot}$ with width $3^{+4.0}_{-1.9}~{}M_{\odot}$ in the PLB model) which may be more easily explained by popular formation channels, which typically do not predict such a narrow feature at $\sim 30-35~{}M_{\odot}$.
+Interestingly, the maximum mass for the PLPLB is $m_{\rm{max},2}=58^{+32}_{-14}~{}M_{\odot}$, and is consistent with the expectation for the onset of the pair–instability supernova (PISN) mass gap (Farmer et al., 2019; Farag et al., 2022), in which case it is possible that the PL2 population also originates from isolated binaries, or is made of first generation dynamically-formed binaries.
+For PL2 from PLPLB, we find a smoothing kernel of $\sim 40~{}M_{\odot}$, revealing a less narrow feature than in the PLB model (found at $32.0^{+1.9}_{-2.0}~{}M_{\odot}$ with width $3.1^{+3.9}_{-1.8}~{}M_{\odot}$ in the PLB model) which may be more easily explained by popular formation channels, which typically do not predict such a narrow feature at $\sim 30-35~{}M_{\odot}$.
 Various formation channels may be able to reproduce a second $>20~{}M_{\odot}$ mass population in the mass distribution which is marginally better represented by a second power law rather than a Gaussian bump.
@@
 The larger width of the distribution compared to PL1 is compelling for a dynamical origin.
-Biscoveanu et al.
-(2022); Heinzel et al.
-(2024b) finds a correlation between the width of the effective spin distribution and redshift.
+Recent work (Biscoveanu et al., 2022; Heinzel et al., 2024a) finds a correlation between the width of the effective spin distribution and redshift.
 We argue that this may be related to what we observe here.
-If we allow for the different sub-populations to vary with redshift independently of each other, we find that at higher redshift the lower mass population rate decreases (as also found by Rinaldi et al.
-2024) while the PL2 rate increases.
-While such redshift evolution may not be physical due to the limited detector horizon especially at lower masses (Heinzel et al., 2024a), this finding may reconcile our and previous works’ observations that the second sub-population drives the spreading of the spin width at higher redshift.
+If we allow for the different sub-populations to vary with redshift independently of each other, we find that at higher redshift the lower mass population rate decreases as also found by (Rinaldi et al., 2024), while the PL2 rate increases.
+While such redshift evolution may not be physical due to the limited detector horizon especially at lower masses (Heinzel et al., 2024b), this finding may reconcile our and previous works’ observations that the second sub-population drives the spreading of the spin width at higher redshift.
 However, note that uncertainties on spins for higher mass systems are typically larger than those from PL1, so it is challenging to draw definitive conclusions.
 Perhaps the most compelling evidence that PL2 may be composed of isolated binaries or dynamically formed first generation binaries is the fact that the PLPLB model PL2 has a maximum mass consistent with PISN mass gap.
-An involve study exploring the redshift and $\chi_{\rm{eff}}$ dependence on the mass distribution is ongoing.
-We note that Godfrey et al.
-(2024) and (Li et al., 2024) also find evidence for 2 distinct sub-populations.
-Specifically, Godfrey et al.
-(2024) finds that the higher mass population which we identify here with PL2 shares characteristics with the likely isolated binary population peaking at $\sim 10~{}M_{\odot}$.
+An involved study exploring the redshift and $\chi_{\rm{eff}}$ dependence on the mass distribution is ongoing.
+We note that (Godfrey et al., 2024) and (Li et al., 2024) also find evidence for 2 distinct sub-populations.
+Specifically, (Godfrey et al., 2024) finds that the higher mass population which we identify here with PL2 shares characteristics with the likely isolated binary population peaking at $\sim 10~{}M_{\odot}$.
 Given all of the arguments and previous works outlined, we argue that PL1 is composed of isolated binaries, while PL2 may mostly arise from 1g+1g binaries from dynamical formation.
-7.2 High-mass $\sim 60-80~{}M_{\odot}$ population
-Although it may be tempting to identify the feature at $\sim 70~{}M_{\odot}$ as the pulsational PISN (PPISN) build-up (Talbot & Thrane, 2018), the bump is located at a mass too high to be at the lower edge of the PISN mass gap, which is expected to be around $45~{}M_{\odot}$ and likely below $56~{}M_{\odot}$ considering reasonable variations in metallicity and uncertainties in the CO reaction rates (Farmer et al., 2019).
+VI.3 High-mass $\sim 60-80~{}M_{\odot}$ population
+Although it may be tempting to identify the feature at $\sim 70~{}M_{\odot}$ as the pulsational PISN (PPISN) build-up (Talbot and Thrane, 2018), the bump is located at a mass too high to be at the lower edge of the PISN mass gap, which is expected to be around $45~{}M_{\odot}$ and likely below $56~{}M_{\odot}$ considering reasonable variations in metallicity and uncertainties in the CO reaction rates (Farmer et al., 2019).
 For the high mass bump, we find a broad distribution on $\chi_{\rm{eff}}$.
@@
 Given these arguments, we argue that a plausible explanation for this feature is from hierarchical mergers of BHs from the $\sim 30-35~{}M_{\odot}$ bump.
-Because we find that the fraction of BBHs in the bump is 1-4% at 90% confidence, assuming these are from 2g BHs in dense star clusters, it is likely that these BHs are born spinning ($\chi_{\rm birth}>0$), but not enough to wash out this feature completely ($\chi_{\rm birth}<0.5$), according to the predictions of Rodriguez et al.
-2019.
-A high mass distribution similar to what found here is also predicted by Kremer et al.
-(2020), with the bump above $\sim 60~{}M_{\odot}$ originating from 2g hierarchical mergers and BHs born from stellar collision.
-8 Conclusions
+Because we find that the fraction of BBHs in the bump is 1-5% at 90% confidence, assuming these are from 2g BHs in dense star clusters, it is likely that these BHs are born spinning ($\chi_{\rm birth}>0$), but not enough to wash out this feature completely ($\chi_{\rm birth}<0.5$), according to the predictions of  (Rodriguez et al., 2019).
+A high mass distribution similar to what found here is also predicted by (Kremer et al., 2020), with the bump above $\sim 60~{}M_{\odot}$ originating from 2g hierarchical mergers and BHs born from stellar collision.
+VII Conclusions
 In this paper we have explored alternative mass models for the BBH mass population.
@@
 We expect that the increasing number of BBH events detected through O4 and beyond will provide the observations required to determine whether the $70~{}M_{\odot}$ sub-population is actually present in the astrophysical BH mass distribution.
-If confirmed, such sub-population would a smoking gun about the dynamical origin of high-mass LVK BBHs, and may represent a new valuable feature to improve cosmological parameters constraints through spectral standard siren measurements (Farr et al., 2019; Ezquiaga & Holz, 2022).
-9 Acknowledgements
+If confirmed, such sub-population would a smoking gun about the dynamical origin of high-mass LVK BBHs, and may represent a new valuable feature to improve cosmological parameters constraints through spectral standard siren measurements (Farr et al., 2019; Ezquiaga and Holz, 2022).
+VIII Acknowledgements
 The authors would like to thank Anarya Ray, Katelyn Breivik, Caitlin Rose and Yeajin Kim for useful comments.